1)
Find the absolute value of the complex number. 12 - 9
2)

Convert the given polar equation to a Cartesian equation. -6 r= 4 cos(O) + sin(0)
Answer #1

1) Given complex number 12-91 absolute value of latbil=fartbr Here a=12, b = -9 112-91 ) - ) (123P+(-73 = V14 144+81 225 = 15